# Learning Video Temporal Dynamics with Cross-Modal Attention for Robust Audio-Visual Speech Recognition

### v1

@@ 3.2 A2V Video Temporal Dynamics Guidance
where is a channel-wise concatenation. The frame order labels are for and for . BCE refers to a binary cross-entropy loss function, and is a binary predictor network, composed of 1D temporal convolution [34] followed by an FC layer. The purpose of temporal convolution is to fuse temporally adjacent features, avoiding ambiguity where a single characteristic (e.g., phoneme) may appear in multiple places within a sequence.
@@ 4.1 Implementation Details
Datasets. We perform our experiments on LRS2 [23] and LRS3 [24], datasets comprising around 224 and 433 hours of audio-visual speech data, respectively, from over 5,000 speakers. Most of our experimental configurations follow [11] and [26], including the noise augmentation and evaluation protocol. We extract noise from MUSAN [25] (babble, music, and natural) and LRS3 (speech) datasets, and partition them into train, validation, and test sets. For training, we sample noise at 0 dB SNR and always add it to the clean speech signal. For evaluation, we use noise from the MUSAN test set, as done in [10, 11, 36], as well as synthesizing the babble noise by randomly mixing 30 audio clips from LRS3, following [5, 7]. We report WER evaluated on noise-perturbed test set with 5 different SNR levels: . The evaluation metric is N-WER [26] (AVG), the average WER across 4 noise categories and 5 SNRs. We also report noise-dominant N-WER (N S), which only considers 3 non-positive SNR levels: .
Model and training description. We adopt AV-HuBERT-Large model [26] as our backbone, which consists of 24 and 9 Transformer [33] layers as the AVSR encoder and decoder, respectively. While more recent AVSR models exist [6, 27, 29], we apply our method to AV-HuBERT for fair comparison with previous works [10, 11, 26, 36] that utilize the same noise-augmenting protocol and pretrained AVSR encoder. As an initialization, we load the pretrained checkpoint from [26], pretrained on noise-augmented LRS3 [24] + VoxCeleb2 [37], and then fine-tune the model for 60K steps on the LRS2 or LRS3 train set. For the first 48K steps, we freeze the AVSR encoder and front-ends of both modalities while training the AVSR decoder, the stacked SA-CA modules, and temporal predictors. We adopt negative log-likelihood for . and are 0.05 and 0.1, respectively. For the binary temporal classifier, we use a kernel size of 3 for the 1D convolutional layer followed by the single FC layer. We set the temporal length as 3, sampling 3 consecutive frames to formulate direction and speed loss functions. The total number of parameters in our whole model is 500M while AV-HuBERT-Large is 477M, implying the stacked SA-CA module only accounts for less than 5%. Our code is implemented upon the fairseq [38] pipeline.
@@ 4.3 Ablation Study
We further examine the performance of visual speech recognition (VSR) to demonstrate the enhancement of visual features without any audio information. Our method, which is accompanied with the video temporal learning, produces a lower 32.5% WER compared to the baseline (33.7%), indicating its better representations of lip movements. However, when video temporal learning is missing or has been misguided by noisy audio, the VSR performance is adversely affected, resulting in 33.2% and 33.3% WER, respectively.
@@
5 Conclusion
In this paper, we have proposed to train the temporal dynamics of video features and employ the cross-modal attention, for the noise-robust AVSR system. Our temporal dynamics learning includes predicting the context order between audio and video frames, the playback direction, and the playback speed of video frames, which significantly enhance the video features based on the temporal dynamics of lip movements. Of our stacked SA-CA module, the cross-modal attention plays a crucial role for correlating each modality to one another. Our methodology achieves the state-of-the-art performance on the LRS2 and LRS3 AVSR benchmarks, particularly when the input audio is perturbed with various noise and SNR level. By extensive ablation studies, we have confirmed the video temporal dynamics learning with cross-modal attention design is essential for improving the noise-robustness of AVSR system.

### v3

@@ 3.2 A2V Video Temporal Dynamics Guidance
where is a channel-wise concatenation. The frame order labels are for and for . BCE refers to a binary cross-entropy loss, and 333For the intuitive loss formulations, the inputs of the binary classifiers in Eqs. 5–3.2 (i.e., and ), are the features that have already undergone 1D temporal convolution in the binary classifier. is a binary predictor network, composed of 1D temporal convolution [34] followed by an FC layer. The purpose of temporal convolution is to fuse temporally adjacent features, avoiding ambiguity where a single characteristic (e.g., phoneme) may appear in multiple places within a sequence.
@@ 4.1 Implementation Details
Datasets. We perform our experiments on the AVSR benchmarks, LRS2 [23] and LRS3 [24]. Most of our experimental configurations follow prior works [11, 26], which include the extraction and addition of noise from MUSAN [25] (babble, music, and natural) and LRS3 (speech) datasets. For training, we sample 0 dB SNR noise and always add it to the clean speech signal for training. We use noise from the MUSAN test set for evaluation as done in [10, 11, 36], as well as the babble noise synthesized from LRS3 [5, 7]. We report WER evaluated on noise-perturbed test set with 5 different SNR levels: . The evaluation metric is N-WER [26] (AVG), the average WER across 4 noise categories and 5 SNRs. We also report noise-dominant N-WER (N S), which only considers 3 non-positive SNR levels.
Model and training description. We adopt AV-HuBERT-Large model [26] as our backbone. As an initialization, we load the pretrained checkpoint from [26], pretrained on noise-augmented LRS3 [24] + VoxCeleb2 [37], and then fine-tune the model for 60K steps on the LRS2 or LRS3 train set. For the first 48K steps, we freeze the AVSR encoder and front-ends of both modalities while training the AVSR decoder, the stacked SA-CA modules, and temporal predictors. We adopt negative log-likelihood for . and are 0.05 and 0.1, respectively. We use a kernel size of 3 for the 1D convolution in the binary temporal classifiers and set the temporal length as 3 in Eqs. 3.2 and 3.2.
@@ 4.3 Ablation Study
We further examine the performance of visual speech recognition (VSR) to demonstrate the enhancement of visual features without any audio information. For the VSR evaluation, the audio features are input as zero vectors, and all the other implementations remain the same. Our method, which is accompanied with the video temporal learning, produces a lower 32.5% WER compared to the baseline (33.7%), indicating its better representations of lip movements. However, when video temporal learning is missing or has been misguided by noisy audio, the VSR performance is adversely affected, resulting in 33.2% and 33.3% WER, respectively.
@@
5 Conclusion
In this paper, we have proposed to train the temporal dynamics of video features and employ the cross-modal attention, for the noise-robust AVSR system. Our temporal dynamics learning includes predicting the context order between two random audio and video frames in each sequence, the playback direction, and the playback speed of video frames. Of our stacked SA-CA module, the cross-modal attention plays a crucial role for correlating each modality to one another. Our methodology achieves the state-of-the-art performance on the LRS2 and LRS3 AVSR benchmarks, particularly when the input audio is perturbed with various noise types and SNR levels. This illustrates the significance of enhancing the video features based on the temporal dynamics of lip movements. By extensive ablation studies about our suggested loss and attention architecture, we have confirmed the video temporal dynamics learning with cross-modal attention design is essential for improving the noise-robustness of AVSR system.
@@ 5 Conclusion
6 Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.20220-00641, 50%] and the National Research Foundation of Korea (NRF) grant funded by MSIT [No. 2021R1A2C1014044, 50%].
Appendix
Appendix A Implementation Details
A.1 Datasets
We perform our experiments on LRS2 [23] and LRS3 [24], datasets comprising around 224 and 433 hours of audio-visual speech data, respectively, from over 5,000 speakers. Most of our experimental configurations follow [11] and [26], including the noise augmentation and evaluation protocol. We extract noise from MUSAN [25] (babble, music, and natural) and LRS3 (speech) datasets, and partition them into train, validation, and test sets. For training, we sample noise at 0 dB SNR and always add it to the clean speech signal. For evaluation, we use noise from the MUSAN test set, as done in [10, 11, 36], as well as synthesizing the babble noise by randomly mixing 30 audio clips from LRS3, following [5, 7].
A.2 Model and Training Description
We adopt AV-HuBERT-Large model [26] as our backbone, which consists of 24 and 9 Transformer [33] blocks as the AVSR encoder and decoder, respectively. This backbone consists of a feature extractor for each modality, where the visual feature extractor is a modified ResNet-18, and the audio feature extractor is a linear projection layer. Extracted features are concatenated to form fusion audio-visual features, which are input to the Transformer encoder. While more recent AVSR models exist [6, 27, 29], we apply our method to AV-HuBERT for fair comparison with previous works [10, 11, 26, 36] that utilize the same noise-augmenting protocol and pretrained AVSR encoder. The hidden dimension of the cross-modal attention is 1024, same as that of the AVSR encoder.
As an initialization, we load the pretrained checkpoint from [26], pretrained on noise-augmented LRS3 [24] + VoxCeleb2 [37], and then fine-tune the model for 60K steps on the LRS2 or LRS3 train set. We observe unstable fine-tuning due to the insertion of untrained cross-modal attention module between pretrained AVSR encoder and front-ends of both modalities. Therefore, we freeze all pretrained modules for the first 48k steps which means only cross-modal attention module and the AVSR decoder are trained during these steps. We adopt negative log-likelihood loss for in a sequence-to-sequence manner with a Transformer decoder. The total number of parameters in our whole model is 500M while AV-HuBERT-Large is 477M, implying the stacked SA-CA module only accounts for less than 5%. Our code is implemented upon the fairseq pipeline.
For the binary temporal classifier, we use a kernel size of 3 for the 1D convolutional layer followed by the single FC layer. We set the temporal length as 3, sampling 3 consecutive frames to formulate direction and speed loss functions. A single frame in our model spans 40 ms, which is shorter than the average English phoneme duration of 100 ms. Thus, we set the kernel size as 3 to match between lip movements and acoustic characteristics such as phonemes, while also minimizing ambiguity from repeated characteristics in a single sequence. Our internal experiments with kernel sizes of 5 and 7 have yielded similar performance to kernel size 3, while leading to slightly increased computational overhead. However, one can consider using a larger kernel size when working with a language that has longer duration phonemes than English.
A.3 Positional Encoding
The conventional Transformer architecture consists of positional encoding (PE) layers prior to self-attention layers to provide contextual information regarding the position of tokens within a sequence. In our framework, which is focused on learning temporal dynamics, we aim to prevent learning redundant information attributable to the PE layer. Consequently, in our cross-modal modules, we apply PE to only the keys and values derived from the other modality streamline, excluding the queries in the CA layer. This implementation ensures that the PE layer does not directly learn temporal dynamics on the video streamline side, where the queries are primarily tasked with computing the temporal losses. For the same reason, SA layers in our cross-modal module are designed without incorporating any PE layers.